Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability -- Deck of Cards -- Help Needed !!!!

Expert replies
by hemanthkumarmn » Wed Jan 30, 2013 12:48 am
Bill has a small deck of 12 playing cards made up of only 2 suits of 6 cards each. Each of the 6 cards within a suit has a different value from 1 to 6; thus, for each value from 1 to 6, there are two cards in the deck with that value. Bill likes to play a game in which he shuffles the deck, turns over 4 cards, and looks for pairs of cards that have the same value. What is the chance that Bill finds at least one pair of cards that have the same value?
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Wed Jan 30, 2013 3:15 am
hemanthkumarmn wrote:Bill has a small deck of 12 playing cards made up of only 2 suits of 6 cards each. Each of the 6 cards within a suit has a different value from 1 to 6; thus, for each value from 1 to 6, there are two cards in the deck with that value. Bill likes to play a game in which he shuffles the deck, turns over 4 cards, and looks for pairs of cards that have the same value. What is the chance that Bill finds at least one pair of cards that have the same value?

A. 8/33
B. 62/165
C. 17/33
D. 103/165
E. 25/33
P(at least 1 pair of the SAME value) = 1 - P(selecting 4 DIFFERENT values).

P(selecting 4 different values):

The first card selected can be ANY VALUE.
P(2nd card is a different value) = 10/11. (Of the 11 cards remaining, any but the mate of the 1st card, leaving 11-1 = 10 good options.)
P(3rd card is a different value) = 8/10. (Of 10 cards remaining, any but the mates of the first 2 cards, leaving 10-2 = 8 good options.)
P(4th card is a different value) = 6/9. (Of the 9 cards remaining, any but the mates of the first 3 cards, leaving 9-3 = 6 good options.)
Since we want all of these events to happen, we multiply the probabilities:
10/11 * 8/10 * 6/9 = 16/33.

Thus:
P(at least 1 pair of the same value) = 1 - 16/33 = 17/33.

The correct answer is C.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by hemanthkumarmn » Wed Jan 30, 2013 3:49 am
GMATGuruNY thanks for your help.. Is there a method to solve this problem without using P(at least 1 pair of the SAME value) = 1 - P(selecting 4 DIFFERENT values). I mean directly finding the at least value..


Hemanth
Join the discussion

by GMATGuruNY » Wed Jan 30, 2013 4:58 am
hemanthkumarmn wrote:GMATGuruNY thanks for your help.. Is there a method to solve this problem without using P(at least 1 pair of the SAME value) = 1 - P(selecting 4 DIFFERENT values). I mean directly finding the at least value..

Hemanth
Favorable outcome 1: Selecting exactly ONE matching pair
Case 1: The first two cards are of the same value, the 3rd and 4th cards are of different values
The 1st card selected can be any of the 12 cards.
P(2nd card is of the same value) = 1/11. (Of the 11 remaining cards, 1 is of the same value as the 1st.)
The next card selected can any of the 10 remaining cards.
P(4th card is of a different value from the 3rd) = 8/9. (Of the 9 remaining cards, 8 are of a different value from the 3rd.)
Since we want both events to happen, we multiply the probabilities:
1/11 * 8/9 = 8/99.

Remaining cases:
Of the 4 cards selected, any combination of 2 could be the matching pair.
Number of combinations of 2 that can be formed from 4 options = 4C2 = (4*3)/(2*1) = 6.
To account for the remaining cases, we multiply by 6:
6 * 8/99 = 16/33.

Favorable outcome 2: Selecting TWO matching pairs
Case 1: The 1st and the 2nd cards match AND the 3rd and the 4th cards match
The first card selected can be any of the 12 cards.
P(2nd card is of the same value) = 1/11. (Of the 11 remaining cards, 1 is of the same value as the 1st.)
The next card selected can any of the 10 remaining cards.
P(4th card is of the same value as the 3rd) = 1/9. (Of the 9 remaining cards, 1 is of the same value as the 3rd.)
Since we want both events to happen, we multiply the probabilities:
1/11 * 1/9 = 1/99.

Remaining cases:
The following ways will yield two matching pairs:
XXYY
XYXY
XYYX
Total ways = 3.
To account for the remaining cases, we multiply by 3:
3 * 1/99 = 1/33.

Since either outcome is favorable, we add the probabilities:
16/33 + 1/33 = 17/33.

The approach suggested in my initial post seems much easier.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by hemanthkumarmn » Wed Jan 30, 2013 5:28 am
GMATGuruNY : True that the earlier solution was lot more simpler, yet out of curiosity wanted to know the alternate.. Anyways thanks a ton..

Hemanth
Join the discussion